OpenStudy (anonymous):

1: Create and provide the dimensions for two similar figures of your choosing. 2: What is the similarity ratio of these figures along with the ratio of their surface area and volume? 3: Show your work, either using the actual volumes or using the formula, that the volume ratio is true. this is what i have: 1: i choose two rectangular prisms. The original dimension are 5x4x10. The new dimensions are 25x16x100. 2: The similarity ratio of theses figures is the new dimensions are the originals squared. The figures are similar and have the same scale factor corresponding sides proportional

OpenStudy (anonymous):

I need help figure out how to write part 3

OpenStudy (anonymous):

Triangular prism has a volume of V=1/2 bhl. If we reduce each dimension by half, then V new =1/2 b/2 h/2 l/2 = 1/16 bhl. Is this what you mean?